OpenStudy (tennistar):

Solve the equation: -4(x + 10) - 6 = -3(x - 2)

OpenStudy (anonymous):

-4x-40-6=-3x+6 -4x+3x=6+6+40 x=-52

OpenStudy (anonymous):

-4(x+10)-6=-3(x-2) -4x-4(10)-6=-3x-3(-2) -4x-40-6=-3x+6 -4x-46=-3x+6 -4x-46+4x=-3x+6+4x -46=x+6 -46-6=x+6-6 x=-52

OpenStudy (tennistar):

thanks

OpenStudy (anonymous):

You are welcome